# Invariant Reduction for Partial Differential Equations. III: Poisson Brackets

### v2

@@ 8.1.2 Reductions of conservation laws
They are $X$-invariant and Poisson commuting. The conservation laws corresponding to $\omega_{3}$ and $\omega_{4}$ are Casimir functions of the Poisson bracket $\{\cdot,\cdot\}_{\nabla}$ on $\mathcal{E}$. Then their reductions are Casimir functions for (8.1). They have the form
@@ 9 Conclusion
The mechanism of invariant reduction provides a natural framework for understanding how symmetry reductions inherit invariant geometric structures. In this paper, we demonstrated that this mechanism also applies to local Hamiltonian operators: under suitable conditions, their reductions give rise to Poisson brackets on the reduced systems. While our focus was on finite-dimensional symmetry reductions, the mechanism is not limited to the finite-dimensional setting. Interpretation of reductions of Hamiltonian operators to infinite-dimensional systems can be based on a direct analog of the formula (4.8). Reductions of the canonical presymplectic structures of cotangent equations can be described in terms of total differential operators, using the $k$-line theorem [26] suitably adapted to the graded-commutative case. Further investigation of these structures and their properties warrants separate study, particularly in the context of multidimensional Hamiltonian systems, including Lagrangian ones.
@@ Acknowledgments
Let us put $\Delta_{\psi}(F)=4D_{x}(\psi)F+2\psi D_{x}(F)$. Then $\Delta_{p}^{*}(p)=4p_{x}p$. The corresponding degree-$1$ symmetry of $\mathcal{E}^{*}$ reads

### v3

@@ 8.1.2 Reductions of conservation laws
They are $X$-invariant and Poisson commuting. The conservation laws corresponding to $\omega_{3}$ and $\omega_{4}$ are Casimir functions of the Poisson bracket $\{\cdot,\cdot\}_{\nabla}$ on $\mathcal{E}$, i.e., they correspond to the zero symmetry. Then their reductions are Casimir functions for (8.1). They have the form
@@ 9 Conclusion
The mechanism of invariant reduction provides a natural framework for understanding how symmetry reductions inherit invariant geometric structures. In this paper, we demonstrated that this mechanism also applies to local Hamiltonian operators: under suitable conditions, their reductions give rise to Poisson brackets on the reduced systems. While our focus was on finite-dimensional symmetry reductions, the mechanism is not limited to the finite-dimensional setting. Interpretation of reductions of Hamiltonian operators to infinite-dimensional systems can be based on a direct analog of the formula (4.8). Reductions of the canonical presymplectic structures of cotangent equations can be described in terms of total differential operators, using the approach via compatibility complexes [27] suitably adapted to the graded-commutative case. Further investigation of these structures and their properties warrants separate study, particularly in the context of multidimensional Hamiltonian systems, including Lagrangian ones.
@@ Acknowledgments
Therefore, $\Delta_{\psi}(F)=4FD_{x}(\psi)+2\psi D_{x}(F)$. Let us put $\Delta_{\psi}=4D_{x}(\psi)+2\psi D_{x}$. Then $\Delta_{p}^{*}(p)=4p_{x}p$. The corresponding degree-$1$ symmetry of $\mathcal{E}^{*}$ reads